OpenStudy (anonymous):

The equation 1/14 (196x+17-10)=______ (20x+12x) has no solution. The equation (4x+24/122)=_____ (x+6/122) Im sure the answer for the second one is 4, (4x+24/122) = 4 (x+6/122)

OpenStudy (mathmale):

I'm having a bit of a problem deciphering (figuring out) what your posted questions are all about. Could you possibly take a screen shot of the original problem and share that screen shot?

OpenStudy (anonymous):

Ok @mathmale

OpenStudy (phi):

for "no solution" you want the "x terms" to be the same on each side for example if you had 2x + 1 = 2x + 2 (notice we have 2x on both sides) if we add -2x to both sides we get 2x - 2x + 1 = 2x - 2x + 2 the 2x-2x is 0 (on both sides) and we have 1=2 and that is not true i.e. there is no x that can make the equation true.

OpenStudy (phi):

(1/14) *196x simplifies to 14 x so you want 14 x on the other side any idea what to multiply 12x by to make it 14x ?

OpenStudy (anonymous):

is it 14/12, since 1/14 (196x+17-10)=(14/12) (20x+12x) makes no solution @phi

OpenStudy (phi):

yes, or simplify 14/12 to 7/6

OpenStudy (mathmale):

Thank you for posting the entire original question. Try to make a habit of doing this in every case.